# Offline Fictitious Self-Play for Competitive Games

### v1

@@ Abstract
Offline Reinforcement Learning (RL) has received significant interest due to its ability to improve policies in previously collected datasets without online interactions. Despite its success in the single-agent setting, offline multi-agent RL remains a challenge, especially in competitive games. Firstly, unaware of the game structure, it is impossible to interact with the opponents and conduct a major learning paradigm, self-play, for competitive games. Secondly, real-world datasets cannot cover all the state and action space in the game, resulting in barriers to identifying Nash equilibrium (NE). To address these issues, this paper introduces Off-FSP, the first practical model-free offline RL algorithm for competitive games. We start by simulating interactions with various opponents by adjusting the weights of the fixed dataset with importance sampling. This technique allows us to learn best responses to different opponents and employ the Offline Self-Play learning framework. In this framework, we further implement Fictitious Self-Play (FSP) to approximate NE. In partially covered real-world datasets, our methods show the potential to approach NE by incorporating any single-agent offline RL method. Experimental results in Leduc Hold’em Poker show that our method significantly improves performances compared with state-of-the-art baselines.
@@ 1 Introduction
Multi-agent reinforcement learning (MARL) provides a powerful learning framework to tackle the problems in multi-agent systems and has been applied to a wide range of domains, such as Go (Silver et al., 2017), strategy games (Vinyals et al., 2019), robotics (Yu et al., 2023), unmanned aerial vehicle (Yun et al., 2022) and network routings (Ye et al., 2015). MARL requires extensive interaction and exploration with accurate and efficient environments, however, this requirement is often not fulfilled in practice, such as football (Kurach et al., 2020), negotiation (Yang et al., 2020), wildlife security protection (Fang et al., 2016), and crowdsourcing (Gerstgrasser et al., 2021). The existing data-driven method, supervised learning, also called behavior cloning, heavily relies on high-quality datasets, which are commonly expensive and suboptimal. Offline MARL offers an excellent alternative to solve these issues by improving policies from previously collected datasets without further interactions (Tseng et al., 2022; Yang et al., 2021).
Recent research in offline MARL has predominantly concentrated on cooperative tasks, where all agents receive a common reward. For cooperative tasks, the goal can be seen as maximizing the cumulative rewards of a joint policy of all agents. Thus, existing single-agent offline RL methods can be easily extended to multi-agent scenarios  (Tseng et al., 2022). However, competitive tasks, often framed as zero-sum games, necessitate a divergent goal orientation.
In the realm of competitive games, or called zero-sum games, exemplified by Texas hold’em poker in studies by (Brown & Sandholm, 2018, 2019), the pursuit of a Nash equilibrium emerges as a prevalent solution concept. Complementary efforts in online setting by (Silver et al., 2017; Ye et al., 2020) focus on the ELO ratio or returns. To reach these goals, These approaches predominantly rely on the self-play paradigm, wherein policies are continuously refined to maximize returns against evolving adversary policies. Nonetheless, in the offline scenario, self-play doesn’t work due to the absence of online interactions, complicating the development of effective offline learning algorithms for competitive games.
In addressing competitive games, recent benchmarks such as AlphaStar Unplugged (Mathieu et al., 2023) and Hokoff (Qu et al., 2023; Wei et al., 2022) have applied single-agent offline RL algorithms to multi-agent settings. However, maximizing the cumulative rewards against static opponents results in an overfitting policy and vulnerable exploitation by more dynamic opponents. Moreover, when datasets are collected by poor policies, such as random policies, it is pointless to learn to defeat these weak opponents, because real-life opponents typically exhibit rational and high-performance behaviors. Li et al. (2022) proposes a model-based offline paradigm for equilibrium finding but requires a strong assumption that datasets fully cover the state-action space of original games. Cui & Du (2022a); Zhong et al. (2022); Cui & Du (2022b) contribute to theoretical insights for a weaker assumption on datasets but only propose algorithms that are feasible in theory. These works encounter challenges in integration with single-agent offline RL methods and do not offer a practical algorithm for addressing real-world problems.
@@
In this paper, we proposed an offline learning framework, called Offline Self-Play (Off-SP), and an offline learning algorithm, Offline Fictitious Self-Play (Off-FSP), for equilibrium finding in zero-sum extensive-form games, bridging the gap between single-agent offline RL and competitive games. To our knowledge, Off-FSP is the first model-free offline algorithm for practical zero-sum game that offers the flexibility to combine with various Offline RL agents. We first propose a technique to approximate interaction with different opponents by re-weighting the datasets with importance sampling. This allows us to learn the best responses against arbitrary opponents with offline RL and derives self-play under the offline setting. By implementing the fictitious self-play on Off-SP, Off-FSP acquires the ability to learn strategies that approximate NE using fixed datasets. We also discuss the challenges of missing data for real-world datasets. The experiments in an imperfect-information two-player zero-sum game, Leduc Hold’em Poker, show that our method outperforms the state-of-the-art baselines and can approach a Nash equilibrium in datasets that fully cover the state-action space of the game. Even with partially covered datasets, our methods show the ability to reduce exploitability in original games.
2 Related Work
In competitive games, also known as zero-sum games, people have developed a series of learning algorithms, where finding Nash equilibrium with self-play is the major learning paradigm for this problem. Fictitious self-play (FSP) (Heinrich et al., 2015) combines fictitious play (FP) (Brown, 1951) with self-play and provably converges to an NE in extensive-form games. Counterfactual regret minimization (CFR) (Bowling et al., 2015) combines regret minimization with self-play, first solving an imperfect-information game of real-world scale. DeepNash (Perolat et al., 2022) extends regularised Nash dynamics with RL and converges to an approximate NE in Stratego. Policy-Space Response Oracles (PSRO) (Lanctot et al., 2017) learn to find NE by iteratively learning best responses to its policies’ population, which can also be seen as a population-based self-play. Some works (Silver et al., 2017; Ye et al., 2020; Yang et al., 2022) of large-scale games aim for a higher winning percentage or returns and try to exploit opponents, not pursuing strict NE, still based on the self-play paradigm.
In the offline learning paradigm, behavior cloning (BC) is the simplest form that directly imitates the sampling strategy of datasets. The performance of BC is limited by the sampling strategy in the dataset, while offline RL can surpass this limitation by maximizing returns. Offline RL faces the challenge of the extrapolation error on out-of-distribution states and actions. Batch-constrained deep q-learning (BCQ) (Fujimoto et al., 2019b) and conservative q-learning (CQL) (Kumar et al., 2020) mitigate this issue by constraining the gap between policy and data distribution. Critic regularized regression (CRR) (Wang et al., 2020) solve this problem by weighted behavior cloning.
In offline competitive scenarios, BC is a viable algorithm, still limited by the quality of sampling strategy in datasets. Mathieu et al. (2023); Qu et al. (2023) directly learn strategies by single-agent offline RL without self-play. This way also requires high-quality behavior strategies in datasets and is easily overfitted to fixed opponents. Cui & Du (2022a); Zhong et al. (2022); Cui & Du (2022b) propose theoretically feasible algorithms. Li et al. (2022) proposed a model-based learning framework, OEF, for extending online equilibrium-finding algorithms in offline scenarios. However, OEF only supports datasets with full coverage of state-action spaces, which is not realistic for real-world problems. Although previous theoretic methods allow for partially covered datasets, they cannot be practically applied to real-life problems. In this paper, Off-SP enables the self-play paradigm in offline MARL of competitive games. Off-FSP implement a practical algorithm to find NE and has the potential to learn on partially covered datasets by integrating any variants of single-agent offline RL algorithms.
@@ 3.1 Extensive-form Game
Extensive-form games are a model of sequential interaction involving $n$ agents. Each player’s goal is to maximize his payoff in the game. At each step $t$ of extensive-form game, only one player observes his respective information states $s^{i}_{t}\in\mathcal{S}^{i}$ and suggests his action $a^{i}_{t}\in\mathcal{A}^{i}(s^{i}_{t})$. $\mathcal{S}^{i}$ is the set of information states of player $i$, and $\mathcal{A}^{i}(s)$ is the set of aviliable action at state $s$. The game’s randomization is defined by an extra player $c$, called chance. Chance follows a fixed randomized strategy that determines the distribution of the next state at its turns. The player function $P:\mathcal{S}\rightarrow\mathcal{N}\cup\{c\}$, with $\mathcal{N}=\{1,\dots,n\}$ denotes the set of players, determines who is to act at a given state.
We assume the games with perfect recall. Each player’s current information state $s_{t}^{i}$ implies the trajectory sequence $(s_{1}^{i},a_{1}^{i},s_{2}^{i},a_{2}^{i},\dots,s_{t}^{i})$, including information states and actions led to $s_{t}^{i}$. In extensive-form games, each player plays following a behavioural strategy that maps all information states to distributions of available actions. The realization-plan (Von Stengel, 1996), $x(s_{t}^{i})=\Pi_{j=1}^{t-1}\pi^{i}(a_{j}^{i}|s_{j}^{i})$, describes the probability of reaching the information state $s_{t}^{i}$ following player $i$’s behavioural strategy, $\pi^{i}$. The strategy profile $\pi=(\pi^{1},\dots,\pi^{n})$, is a joint of all player’s behavioural strategy. $\pi^{-i}$ denotes the strategy profile of all players excepts $i$. The payoff of player $i$ is denoted by $R^{i}(\pi^{i},\pi^{-i})\in\mathbb{R}$, and $\sum_{i}R^{i}=0$ for zero-sum games. Given a fixed $\pi^{-i}$, best response $\textsc{BR}(\pi^{-i})$ is the strategy that has the highest payoff. A Nash equilibrium (NE) is a strategy profile that any strategy $\pi^{i}$ in this profile is a best response to the opponent’s profile $\pi^{-i}$. The $\epsilon$-best response ($\epsilon$-BR) and $\epsilon$-Nash equlibirium ($\epsilon$-NE) are approximations to above defination. $\epsilon$-BR is suboptimal by no more than $\epsilon$ compared with BR. Similarly, $\epsilon$-NE is a profile of $\epsilon$-BR. NashConv (Timbers et al., 2022), also called exploitability in two-player zero-sum games, evaluates the distance from $\pi$ to an NE, defined as $\sum_{i}R^{i}(\textsc{BR}(\pi^{-i}),\pi^{-i})-R^{i}(\pi^{i},\pi^{-i})$.
Normal-form game is a simplified game description in game theory. Normal form represents the game by way of a matrix. An extensive-form game induces an equivalent normal-form game. In Sections 4.4 and 5.1, we discuss with toy examples of a normal-form game, rock-paper-scissors, and show the experimental results.
Algorithm 1 Fictitious Self-Play
Function FictitiousSelfPlay()
      Initialize strategy $\pi_{0}$.  for $k=1$ to $K$ do
           $\mathcal{D}_{k}\leftarrow$ GenerateData($\pi_{k-1}$) foreach player $i\in\mathcal{N}$ do
                $\beta^{i}_{k}\leftarrow$ LearnBestResponse($\mathcal{D}_{k},\pi^{-i}_{k-1}$)
          $\pi_{k}\leftarrow$ UpdateAverageStrategy($\pi^{i}_{k-1},\beta^{1}_{k},\dots,\beta^{n}_{k}$)
     return average strategy $\pi_{K}$
@@ 3.2 Fictitious Self-Play
Fictitious Play (Brown, 1951) is a game-theoretic model for learning from self-play. Fictitious players iteratively compute the best responses to opponents’ average behavioural strategy and update their set of strategies. The average strategies of fictitious players are proven to converge to Nash equilibria in two-player zero-sum and many-player potential games. Generalised weakened fictitious play (Leslie & Collins, 2006) is an improvement of fictitious play that enables approximate best responses and perturbed average strategy updates. Generalised weakened fictitious play also provably converges to Nash equilibria in certain game classes and is suitable for machine learning framework.
Original fictitious Play is defined in normal-form games, which requires exponential computations for extensive-form games. Heinrich et al. (2015) introduce Fictitious Self-Play (FSP) that extend fictitious play and generalised weakened fictitious play into the extensive-form game with linear time and space complexity. We briefly describe FSP at  Algorithm 1. In extensive-form games, the average strategies $\pi^{i}_{k}$ are updated by the realization-equivalence theorem. At the $k$-th iteration of FSP, the average behavioural strategy of player $i$ is
$\begin{split}\forall s,a:~{}&\pi^{i}_{k}(a|s)=(1-\lambda)\pi^{i}_{k-1}(a|s)+\lambda\beta^{i}_{k}(a|s),\\ &\lambda=\frac{\alpha_{k}x_{\beta^{i}_{k}}(s)}{(1-\alpha_{k})x_{\pi^{i}_{k-1}}(s)+\alpha_{k}x_{\beta^{i}_{k}}(s)},\end{split}$ (1)
where $\beta^{i}_{k}\in\epsilon_{k}\textsc{-BR}(\pi_{k-1}^{-i})$ is a best-response to opponent $\pi_{k-1}^{-i}$ and $\alpha_{k}$ is parameter to mix the best-response $\beta^{i}_{k}$ and previous average strategy $\pi^{i}_{k-1}$. Strategy $\pi^{i}_{k}$ is equivalent to choosing either strategy $\pi^{i}_{k-1}$ or $\beta^{i}_{k}$ before the beginning of each game, with probabilities $1-\alpha_{k}$ and $\alpha_{k}$ respectively. When $\alpha_{k}\rightarrow 0$ and $\epsilon_{k}\rightarrow 0$ as $k\rightarrow 0$, and $\sum_{k=1}^{\infty}\alpha_{t}=\infty$ for all $n$ players, the average strategy profile $\pi_{k}=(\pi^{1}_{k},\dots,\pi^{n}_{k})$ converge to an NE. In a common setting with $\alpha_{k}=\frac{1}{k}$, $\pi_{K}^{i}$ is equivalent to choosing all strategy $\beta^{i}_{1},\dots,\beta^{i}_{K}$ with the probability of $\frac{1}{K}$. The update of average behavioural strategy requires a traverse for all information state $s^{i}\in\mathcal{S}^{i}$, which is computationally expensive for large games and machine learning-based frameworks. FSP restored all the data of strategies $\beta^{i}_{1},\dots,\beta^{i}_{k}$ and trained an approximate average strategy with supervised learning.
@@ 3.3 Offline Reinforcement Learning
(a) Off-FSP and Off-SP.
(b) Differences between FSP and Off-FSP.
Figure 1: Illustration of Off-FSP, Off-SP and three essential functions. The green box in (a) is the Off-SP.
@@
Reinforcement learning (RL) agents aim to maximize the expected cumulative discounted reward in a Markov decision process (MDP). MDP is denoted as a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{T},\rho_{0},r,\gamma,T)$. $\mathcal{S},\mathcal{A}$ represent the state and action spaces. $\mathcal{T}(s_{t+1}|s_{t},a_{t}),r(s_{t},a_{t})$ represent the dynamics and reward function. $\rho_{0}$ is the distribution of initial state $s_{1}\sim\rho_{0}$. $\gamma\in[0,1]$ is the discount factor. Many reinforcement learning algorithms learn from transition tuples, $(s_{t},a_{t},r_{t},s_{t+1})$, where $s_{t},a_{t},r_{t}$ are the state, action and reward at time $t$, and $s_{t+1}$ is the next state after agent take action $a_{t}$ at state $s_{t}$. The expected cumulative discounted reward following policy $\pi$ can be formalized as the action-value function $Q(s_{t},a_{t})=\mathbb{E}_{\pi}[\sum^{\infty}_{i=t}\gamma^{i-t}r(s_{i},a_{i})]$. Model-based RL methods learn the dynamic models while model-free methods directly learn the Q function to state-action pairs. Given the fixed opponent $\pi_{-i}$ and an extensive form game, the game of player $i$ can be defined as a MDP $\mathcal{M}(\pi_{-i})$ (Silver & Veness, 2010; Greenwald et al., 2013). Player $i$’s information state defines the state, and his payoff function defines the reward. An $\epsilon$-optimal policy of the MDP, $\mathcal{M}(\pi^{-i})$, is also the $\epsilon$-BR to the strategy $\pi^{-i}$.
@@
Offline RL algorithm breaks the assumption that the agent can interact with the environment. The offline RL algorithm learns to maximize the expected cumulative reward based on a fixed dataset. The dataset is sampled following a behavior policy $\pi_{b}(a|s)$ in MDP $\mathcal{M}$. $\hat{\pi}_{b}(a|s):=\frac{\sum_{s,a\in\mathcal{D}^{i}}\mathbbm{1}[s=s,a=a]}{\sum_{s\in\mathcal{D}^{i}}\mathbbm{1}[s=s]}$ denote the empirical behavior policy, at all state $s\in\mathcal{D}^{i}$. In offline RL, the Q-function is trained only on actions sampled from offline datasets, which cannot cover all state and action spaces. Therefore, the Q-function may be erroneously overestimated at out-of-distribution (OOD) actions, which is called extrapolation error (Fujimoto et al., 2019b). Recent offline RL methods encourage the policy to learn on the support of training data or employ weighted behavior cloning to mitigate this error.

### v2

@@ Abstract
Offline Reinforcement Learning (RL) enables policy improvement from fixed datasets without online interactions, making it highly suitable for real-world applications lacking efficient simulators. Despite its success in the single-agent setting, offline multi-agent RL remains a challenge, especially in competitive games. Firstly, unaware of the game structure, it is impossible to interact with the opponents and conduct a major learning paradigm, self-play, for competitive games. Secondly, real-world datasets cannot cover all the state and action space in the game, resulting in barriers to identifying Nash equilibrium (NE). To address these issues, this paper introduces Off-FSP, the first practical model-free offline RL algorithm for competitive games. We start by simulating interactions with various opponents by adjusting the weights of the fixed dataset with importance sampling. This technique allows us to learn the best responses to different opponents and employ the Offline Self-Play learning framework. To overcome the challenge of partial coverage, we combine the single-agent offline RL method with Fictitious Self-Play (FSP) to approximate NE by constraining the approximate best responses away from out-of-distribution actions. Experiments on matrix games, extensive-form poker, and board games demonstrate that Off-FSP achieves significantly lower exploitability than state-of-the-art baselines. Finally, we validate Off-FSP on a real-world human-robot competitive task, demonstrating its potential for solving complex, hard-to-simulate real-world problems.
@@ 1 Introduction
Multi-agent reinforcement learning (MARL) provides a powerful learning framework to tackle the problems in multi-agent systems and has been applied to a wide range of domains, such as Go (Silver et al. 2017), strategy games (Vinyals et al. 2019), robotics (Yu et al. 2023), unmanned aerial vehicle (Yun et al. 2022), and network routings (Ye, Zhang, and Yang 2015). MARL typically relies on extensive interaction and exploration with accurate and efficient environments, hindering its application in the real world. In many real-world multi-agent problems, such as football (Kurach et al. 2020), negotiation (Yang, Chen, and Narasimhan 2020), and crowdsourcing (Gerstgrasser, Trivedi, and Parkes 2021), the absence of reliable simulators makes it expensive and inefficient to interact with the environment and collect new data for training MARL agents. Moreover, in special problems, such as wildlife protection (Fang et al. 2016), learning MARL agents online by trial and error is unsafe, which could lead to danger for patrols or wildlife. In these cases, learning from existing data can be extremely valuable as it does not require additional sampling. Offline MARL offers an excellent alternative to solve these issues by improving policies from previously collected datasets without further interactions (Tseng et al. 2022; Yang et al. 2021).
Figure 1: Comparison of offline self-play with other learning paradigms for real-world competitive tasks. Online self-play requires a simulator and suffers from the sim-to-real gap. Behavior cloning needs costly expert data. Naive offline RL optimizes with incorrect objectives. In contrast, Off-SP learns from inexpensive, low-quality data. See Section˜5.2 for experiments on the illustrated real-world game.
In competitive multi-agent games, often framed as zero-sum games, the objective is to find a Nash equilibrium (NE) (Kreps 2018) or maximizing the ELO ratio (Silver et al. 2017; Ye et al. 2020), which is divergent from the goal of maximizing cumulative rewards in single-agent situations. Exemplified by studies of Texas hold’em poker (Brown and Sandholm 2018, 2019), the pursuit of a Nash equilibrium enables agents to learn robust policies against various opponents, enhancing their performance in competitive environments. To achieve the goals, online approaches predominantly rely on the self-play paradigm (Zhang et al. 2024), wherein policies are continuously refined to maximize returns against evolving adversary policies. However, in the offline scenario, the challenge arises from the absence of online interactions with evolving opponents, complicating the development of self-play paradigms. Recent benchmarks of offline competitive games, such as AlphaStar Unplugged (Mathieu et al. 2023) and Hokoff (Qu et al. 2023; Wei et al. 2022), have directly applied single-agent offline RL algorithms with no self-play. However, maximizing the cumulative rewards against static opponents results in an overfitting policy and vulnerable exploitation by more dynamic opponents. Moreover, when datasets are collected by poor policies, such as random policies, it is pointless to learn to defeat these weak opponents, because real-life opponents typically exhibit rational and high-performance behaviors.
The reliance on high-quality or high-coverage datasets, which are commonly expensive and suboptimal in the real world, is another challenge for offline learning. The existing data-driven method, supervised learning, also called behavior cloning, ignores the objective of online learning and only imitates the sampling policy of datasets. Consequently, this method performs poorly when dealing with non-expert datasets. Many recent works are learning to improve policies toward the NE, but they demand high state-action coverage within the dataset. Li et al. (2022) proposes a model-based offline paradigm for equilibrium finding, but it requires a strong assumption that datasets fully cover the state-action space of the original games. Cui and Du (2022b); Zhong et al. (2022); Cui and Du (2022a) contribute to theoretical insights for a weaker assumption on datasets but only propose theoretically feasible algorithms. Existing work lacks a practical method applicable to non-expert and partially covered real-world datasets. Figure˜1 compares our paradigm with existing works, highlighting the advantage of Off-SP on solving real-world competitive problems. Detailed comparisons with related works are provided in the Appendix˜B.
@@
In this paper, we propose an offline learning framework, called Offline Self-Play (Off-SP), and an offline learning algorithm, Offline Fictitious Self-Play (Off-FSP), for equilibrium finding in zero-sum extensive-form games, bridging the gap between single-agent offline RL and competitive games. To our knowledge, Off-FSP is the first model-free offline algorithm for practical zero-sum games that offers the flexibility to combine with various Offline RL agents and improve policies on non-expert real-world datasets. We first propose a technique to approximate interaction with different opponents by re-weighting the datasets with importance sampling. This allows us to learn the approximate best responses against arbitrary opponents with offline RL and derive a self-play paradigm under the offline setting. Also, for partially covered and non-expert datasets, we use a surrogate loss, NashConv, to measure the distance to NE, rather than finding the exact NE. Off-FSP combines single-agent offline reinforcement learning methods with fictitious self-play to learn approximate best responses iteratively and minimize the NashConv. We validate our approach across matrix-form and extensive-form zero-sum games, demonstrating consistently low exploitability and superior performance over existing offline RL baselines under partially covered datasets. Notably, we further apply our method to a real-world human-robot competitive task, showcasing its potential in addressing hard-to-simulate decision-making problems.
@@ 2.1 Extensive-form Game
Extensive-form games are a model of sequential interaction involving $n$ agents. Each player’s goal is to maximize his payoff in the game. At each step $t$ of extensive-form game, only one player observes his respective information states $s^{i}_{t}\in\mathcal{S}^{i}$ and suggests his action $a^{i}_{t}\in\mathcal{A}^{i}(s^{i}_{t})$. $\mathcal{S}^{i}$ is the set of information states of player $i$, and $\mathcal{A}^{i}(s)$ is the set of available action at state $s$. The player function $\mathcal{P}:\mathcal{S}\rightarrow\mathcal{N}$, with $\mathcal{N}=\{1,\dots,n\}$ denotes the set of players, determines the player to act.
In extensive-form games, each player plays following a policy $\pi^{i}:\mathcal{S}^{i}\rightarrow\Delta(\mathcal{A}^{i})$ that maps information states to distributions of actions. The realization-plan (Von Stengel 1996), $x(s_{t}^{i})=\Pi_{j=1}^{t-1}\pi^{i}(a_{j}^{i}|s_{j}^{i})$, describes the probability of reaching the information state $s_{t}^{i}$ following player $i$’s policy, $\pi^{i}$. The strategy profile $\pi=(\pi^{1},\dots,\pi^{n})$, is a joint of all player’s policy. $\pi^{-i}$ denotes the strategy profile of all players excepts $i$. The payoff of player $i$ is denoted by $R^{i}(\pi^{i},\pi^{-i})\in\mathbb{R}$, and $\sum_{i}R^{i}=0$ for zero-sum games. Given a fixed $\pi^{-i}$, best response $\textsc{BR}(\pi^{-i})$ is the policy with the highest payoff. A Nash equilibrium (NE) is a strategy profile that any policy $\pi^{i}$ in this profile is a BR to the opponent’s profile $\pi^{-i}$. The $\epsilon$-best response ($\epsilon$-BR) and $\epsilon$-Nash equlibirium ($\epsilon$-NE) are approximations to the above definition. $\epsilon$-BR is suboptimal by no more than $\epsilon$ compared with BR. Similarly, $\epsilon$-NE is a profile of $\epsilon$-BR. NashConv (Timbers et al. 2022), also called exploitability in two-player zero-sum games, evaluates the distance from $\pi$ to an NE, defined as $\sum_{i}R^{i}(\textsc{BR}(\pi^{-i}),\pi^{-i})-R^{i}(\pi^{i},\pi^{-i})$.
@@ 2.2 Fictitious Self-Play
Fictitious Self-Play (FSP)(Heinrich, Lanctot, and Silver 2015) is a game-theoretic model that iteratively computes the best responses to opponents’ average policy and updates their set of policies. We briefly describe FSP at  Appendix˜A. In extensive-form games, the average policies $\pi^{i}_{k}$ are updated by the realization-equivalence theorem. At the $k$-th iteration of FSP, the average policy $\pi^{i}_{k}$ of player $i$ is
$\begin{split}\forall s,a:~&\pi^{i}_{k}(a|s)=(1-\lambda)\pi^{i}_{k-1}(a|s)+\lambda\beta^{i}_{k}(a|s),\\ \lambda&=\frac{\alpha_{k}x_{\beta^{i}_{k}}(s)}{(1-\alpha_{k})x_{\pi^{i}_{k-1}}(s)+\alpha_{k}x_{\beta^{i}_{k}}(s)},\end{split}$ (1)
where $\beta^{i}_{k}\in\epsilon_{k}\textsc{-BR}(\pi_{k-1}^{-i})$ is a best-response to opponent $\pi_{k-1}^{-i}$ and $\alpha_{k}$ is the mixing parameter. Policy $\pi^{i}_{k}$ is equivalent to choosing either policy $\pi^{i}_{k-1}$ or $\beta^{i}_{k}$ before the beginning of each game, with probabilities $1-\alpha_{k}$ and $\alpha_{k}$ respectively. A standard choice is $\alpha_{k}=\frac{1}{k}$.
Figure 2: Example Datasets of RPS. Numbers in the grids show the probability density of different samples. The red dashed boxes indicate the probability of different actions for corresponding behavioural policies.
Figure 3: Results of RPS. The prefix of D1- and D2- are refering to restuls on the first and second datasets respectively.
@@ 2.3 Offline Reinforcement Learning
Reinforcement learning (RL) agents aim to maximize the expected cumulative discounted reward in a Markov decision process (MDP). MDP is denoted as a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{T},\rho_{0},r,\gamma,T)$. $\mathcal{S},\mathcal{A}$ represent the state and action spaces. $\mathcal{T}(s_{t+1}|s_{t},a_{t}),r(s_{t},a_{t})$ represent the dynamics and reward function. $\rho_{0}$ is the distribution of initial state $s_{1}\sim\rho_{0}$. $\gamma\in[0,1]$ is the discount factor. The expected cumulative discounted reward following policy $\pi$ can be formalized as the action-value function $Q(s_{t},a_{t})=\mathbb{E}_{\pi}[\sum^{\infty}_{i=t}\gamma^{i-t}r(s_{i},a_{i})]$. Given the fixed opponent $\pi_{-i}$ and an extensive form game, the game of player $i$ can be defined as a MDP $\mathcal{M}(\pi_{-i})$ (Silver and Veness 2010; Greenwald et al. 2013). An $\epsilon$-optimal policy of the MDP, $\mathcal{M}(\pi^{-i})$, is also the $\epsilon$-BR to the policy $\pi^{-i}$.
@@
Offline RL algorithm breaks the assumption that the agent can interact with the environment. The offline RL algorithm learns to maximize the expected cumulative reward based on a fixed dataset. The dataset is sampled following a behavior policy $\pi_{b}(a|s)$ in MDP $\mathcal{M}$. $\hat{\pi}_{b}(a|s):=\frac{\sum_{s,a\in\mathcal{D}^{i}}\mathds{1}[s=s,a=a]}{\sum_{s\in\mathcal{D}^{i}}\mathds{1}[s=s]}$ denote the empirical behavior policy, at all state $s\in\mathcal{D}^{i}$. The Q-function may be erroneously overestimated at out-of-distribution (O. O. D.) actions, which is called extrapolation error (Fujimoto, Meger, and Precup 2019). Recent offline RL methods encourage the policy to learn on the support of training data or employ weighted behavior cloning to mitigate this error.
3 The Motivating Example: Rock-Paper-Scissors
We start by analysing the well-known game, Rock-Paper-Scissors (RPS), to illustrate the challenges of learning in offline datasets of competitive games. In this game, two players can choose one of three actions: Rock, Paper, or Scissors. The payoffs are defined as follows: Rock beats Scissors, Scissors beats Paper, and Paper beats Rock. The game has a unique Nash equilibrium (NE) where both players play each action with equal probability of $\frac{1}{3}$. In Figure˜3, we show two datasets of RPS. The first dataset is a fully covered dataset sampled from a non-expert policy, $P(\text{R,P,S})=(0.6,0.2,0.2)$. The second dataset is a partially covered dataset sampled from a modified asymmetric RPS game, where the second player has a new action, Rock2, with the same payoff as Rock.
In the first non-expert dataset, neither Behavioral Cloning (BC) nor naive offline RL can find the NE, which highlights the importance of the self-play paradigm. BC mimics the distribution of the dataset by supervised learning. Its policy, $P_{\text{BC}}(\text{R,P,S})=(0.6,0.2,0.2)$, is suboptimal and beaten by the policy of playing Paper only. As the solution of  Mathieu et al. (2023); Qu et al. (2023), single-agent offline RL maximizes the payoff under the fixed opponent in the dataset, and its policy, $P_{RL}(\text{R,P,S})=(0,1,0)$, is also easy to be exploited by another policies. Under the self-play paradigm, the policy learns to play against a dynamic opponent, which is more robust and can approximate the NE. Figure˜3 compares the performance of our method, Off-FSP-CQL, with the baselines, BC, Off-FSP-DQN and online Fictitious Play (FP) (Brown 1951).
Corresponding to the challenge of learning in partially covered datasets, the second dataset in Figure˜3 leaves the payoff of Rock2 not fully observed. In such datasets, finding the accurate NE of the original game is impossible , so the goal of learning is to minimize the exploitability and NashConv, i.e., the distance to NE. In the dataset, the only observed sample of Rock2 is $(\text{Scissors},\text{Rock2})$. In player $2$’s perspective, the Rock2 action gets a payoff of $1$ all the time, so choosing it with a probability of $1$ is the best choice. Nevertheless, the policy has the highest exploitability in the original game, as player $1$ possesses a best response by selecting scissors, resulting in a payoff of $1$ for player $1$. Offline RL algorithms, such as Conservative Q-Learning(CQL) (Kumar et al. 2020), mitigate this problem by punishing the agent for learning OOD actions, including unseen actions and undersampled actions. By combining self-play and offline RL, Off-FSP ignores the samples with large uncertainty and learns robust policies with low exploitability. As shown in  Figure˜3, Off-FSP without offline RL (Off-FSP-DQN) converges to high exploitability, while Off-FSP with CQL can reduce the exploitability to a low level. We also show the learning curve of an online algorithm, FP, which shows the convergence speed of Off-FSP is comparable to online algorithms. Further explanations of the results are given in the Appendix˜E.
((a)) Offline Self-Play framework.
((b)) Pipeline of Off-FSP.
((c)) Differences between FSP and Off-FSP.
Figure 4: Illustration of Off-SP, Off-FSP and three essential steps. The green box in (b) is Off-SP.